# Misspecification uncertainties in near-deterministic regression

### v2

@@ IV.2 Test errors in medium-dimensional linear regression
We now look at higher/medium dimensional problems, $P=10-100$, where direct minimization of ${\mathcal{G}}$ is not practical. The simulation engines (see supplementary material) were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models. A noise term that simulates the presence of variability orthogonal to the chosen feature set was also used. In all cases, independent test and training data were generated, typically with a 10:90 test:train split. However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered. Additional examples are provided in the supplementary material and appendix G.
IV.2.1 POPS-ensemble prediction of test errors
For each test point, it is simple to generate an ‘envelope’, or max/min range, from the POPS-ensemble. We find this envelope provides a highly informative bound on test errors for large $N/P$. The top left of figure 3 is a parity plot for the minimum loss model, showing around 93% coverage, or 7% envelope violation. The violation rate drops with $N/P$, as shown on the bottom of figure 3 for a range of $P$ and $N/P$ and two simulation engines.
Whilst the envelope bounds are highly informative, it is also desirable to predict test error distributions. As discussed above, the limitation of our POPS-ensemble is that the weighting of ensemble members is unlikely to be optimal, whilst the max/min range can be expected to be robust. As a result, a conservative estimate of the test error can be generated by uniformly resampling the max/min bounds for each test point to generate predicted error histograms. As shown in the top (right) of figure 3, this prediction from the POPS-ensemble results in excellent agreement with the test error ensemble, as confirmed across multiple engines in appendix B. Whilst a single (re)sample per test point is sufficient, it is possible to effectively bootstrap error histograms through repeated resampling.
IV.2.2 Resampled POPS (rPOPS) ensemble
The parameter vectors of the POPS-ensemble across all simulation engines studied are observed to be dense across their simplex, or hypercube of the max/min range per parameter. Inspired by the success of the uniform resampling approach used to produce test error predictions of the POPS ensemble in figure 3, we have also investigated the effect of uniformly resampling this ‘POPS-simplex’, a process we name as ‘resampled’ POPS, or rPOPS. We emphasize that the rPOPS approach makes a strong assumption that all parameter vectors inside the POPS- simplex are close to valid POPS-ensemble members. We thus include these results as a motivation for further study; in particular, whilst the $\mathcal{O}(N)$ cost of the POPS-ensemble remains highly managable, the cost of the rPOPS-ensemble can be tailored through the number of resamples, offering large efficiencies as $N/P\to\infty$ when producing envelopes and test error distributions.
As shown in figure 3, the rPOPS results are excellent, giving tight agreement with test error distributions and only slightly more conservative errors bounds, a feature which is again robust across simulation engines (appendix G). A detailed investigation of the POPS-simplex and theoretical analysis of the rPOPS-ensemble is left as a topic for future study. We conclude with an application to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
@@ IV.2.2 Resampled POPS (rPOPS) ensemble
Figure 3: Top: test error predictions for a misspecified linear surrogate model (P=20) on a cubic simulation engine. Left: parity plot for the minimum loss model, along with envelope bounds from the POPS-ensemble and rPOPS-ensemble (see text), showing 93-98% coverage at $N/P=750$. Right: Test error histogram, along with predicted test errors from the POPS- and rPOPS-ensemble, as explained in the text. Below: POPS-ensemble envelope violation rates with $P$ and $N/P$ for two simulation engines.
IV.3 Interatomic potentials for material simulations
Figure 4: Characteristics of the ensemble method for the application to interatomic potentials. Left: Parity plot for energy and force predictions for $N/P=50$. The errorbars reflect the min/max envelope of the POPS-ensemble. Right: cumulative distribution of the ratio of the residual of the minimum-loss model to the envelope width (max/min range per test point) for different conditions.
@@ IV.3 Interatomic potentials for material simulations
In the following, training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022). 10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components). Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), which can be expressed as a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq. 23. The properties of the ensemble are then evaluated on a complementary set of 1,000 testing points. The random test/train partition is repeated 10 times for each value of $N/P$.
The results show that the envelope violation on test data varies between $40\%$ and $10\%$ for $N/P=1$ and $50$, respectively, in qualitative agreement with the results shown in figure 3, demonstrating that worst-case model errors are increasingly well captured by the ensemble envelope. Indeed, the scaling is actually superior to the medium test examples. As seen in these examples, the envelope induced by the model ensemble yields a useful estimate of the actual errors, e.g., as incurred by the ordinary minimum-loss model, as reported in figure 4. For example, the mean ratio of the residual of the minimum-loss model to the envelope width (max/min range) is around $0.45$ at $N/P=1$, i.e, the envelope width overestimates the actual minimum-loss residual only by a factor of around 2 on average. This figure drops to about $0.25$ at $N/P=50$, following the broadening of the envelopes that occurs as rare ”outliers” that are particularly difficult to capture within the model class are gradually added to the training set. This was empirically verified by sampling ensembles containing only $P$ models, independently of the size of the training set: in this case, the mean ratio of the residual of the minimum-loss model to the envelope width remains around $0.45$ (c.f., figure 4). This numerical experiment indicates that the width of the model envelope predictions provides reliable worst-case confidence intervals as well as representative (albeit somewhat conservative) estimates of the actual errors at testing points.
V Conclusion
@@
stIn spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation. In this work, we show that the important near-deterministic underparameterized regime is amenable to approximation by a formally simple discrete ensemble ansatz that exploits the concentration of the cross-entropy minimizing ensemble onto pointwise optimal parameter sets. For linear regression problems, this ensemble can be obtained extremely efficiently using rank-one perturbations of the expected loss solution. POPS-constrained loss-minimizing ensembles are shown to produce excellent parameter distributions, worst-case bounds on model predictions, and approximations of the pointwise errors.
Software and Data
Will be released after peer review
@@ Acknowledgements
This work was initiated during a visit at the Institute for Pure and Applied Mathematics (IPAM) at the University of California in Los Angeles (UCLA). Their hospitality is graciously acknowledged. TDS gratefully acknowledges support from ANR grant ANR-19-CE46-0006-1, IDRIS allocations A0120913455 and an Emergence@INP grant from the CNRS. D.P. was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20220063DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).
@@ Appendix A PAC-Bayes concentration inequalities for regression
Determining the value of the upper bound for $C_{0}[\pi_{0}]$ in the PAC-Bayes concentration inequality (7) requires certain assumptions on the distribution of the error between the empirical and true expected loss (Germain et al., 2016). Following (Germain et al., 2016), we assume that this error distribution is unbounded but sub-normal, i.e. all error fluctuations are bounded by a normal distribution of variance $s^{2}$, giving $C_{0}[\pi_{0}]<s^{2}/2$. We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion.
@@ Appendix B Model selection criteria for deterministic models
We note that the value of the PAC-Bayes loss bound (7) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in (Germain et al., 2016). In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected. However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed. Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{E}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (25)
@@ Appendix C Summary of Laplace’s method
Appendix D POPS-constrained variational minimization of ensemble ansatz

### v3

@@ III.2 Test errors in medium-dimensional linear regression
We now consider higher/medium dimensional problems, $P=10-50$, where numerical minimization of ${\mathcal{G}}_{E}$ is not possible. The simulation engines were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models. A noise term that simulates the presence of variability orthogonal to the chosen feature set was also used. In all cases, independent test and training data were generated, typically with a 10:90 test:train split. However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered Swinburne and Perez (2024).
As shown, the $\mathcal{O}(P)$ hypercube ansatz $\pi^{*}_{\mathcal{H}}$ provides an excellent prediction and bounding, of the test error distribution whilst the $\mathcal{O}(N)$ ensemble $\pi^{*}_{E}$ underestimates moments of the error distribution and the max/min envelope bound. Importantly, the envelope violation rate of the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ drops with $N/P$, as shown in figure 2 for a range of $P$ and $N/P$. The simultaneous prediction and bounding of test errors is highly valuable for surrogate model deployment, allowing not only the prediction of expected errors but a robust assessment of worst case error scenarios. We conclude with an application of the hypercube ansatz to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
@@
Figure 2: Test errors of a misspecified linear surrogate model on a cubic simulation engine. Left: test error histogram at $P=20$, $N/P=100$ for the minimum loss model (black) and predictions from the POPS-ensemble $\pi^{*}_{E}$ (orange) and POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz. MAE: mean absolute error relative to the minimum loss solution. EV: envelope violation, points lying outside of the max/min bound. Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $P$ and $N/P$.
III.3 High-dimensional models for atomic simulations
Figure 3: Application to interatomic potentials. Left: test error histogram at $P=1596$, $N/P=81$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz. MAE: mean absolute error relative to the minimum loss solution. EV: envelope violation, points lying outside of the max/min bound. Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$.
@@ III.3 High-dimensional models for atomic simulations
In the following, a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}$ was used to predict per-atom contributions. Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq. 20, up to a maximal value of $N=129,853$. The properties of the ensemble are then evaluated on a complementary set of 10,000 hold-out testing points. Error analysis is carried out in the context of the POPS-hypercube ansatz. Even on the full dataset, the analysis requires on the order of 20 minutes on 1 CPU core. It is therefore extremely lightweight in addition to being easily parallelizable, enabling its use for extremely large datasets.
The results show that the envelope violation on test data varies between $2\%$ and $0.2\%$ for $N/P=0.06$ and $33$, respectively, in qualitative agreement with the results shown in figure 2. This demonstrates that worst-case model errors are extremely well captured by the POPS-hypercube ensemble. As reported in the left panel of figure 3, these excellent bounds do not trivially result from the resampled error distribution being overly pessimistic, as the distribution of errors induced by a uniform parameter resampling within the POPS-hypercube provides an excellent predictor of the actual error distribution with respect to the minimum-loss model. The MAE estimated from the resampled models exceeds the measured value by only 23%, while providing a very good description of the overall error distribution.
The information generated by this simple and efficient approach will be invaluable to estimate the uncertainties on a range of different physical quantities inferred from atomistic simulations.
IV Conclusion
@@
In spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation. In this work, we show that the important near-deterministic underparameterized regime is amenable to approximation by two formally simple ansatz that exploits the concentration of the cross-entropy minimizing ensemble onto pointwise optimal parameter sets (POPS). For linear regression problems, this ensemble can be obtained extremely efficiently using rank-one perturbations of the expected loss solution, from which our optimally weighted POPS-ensemble or POPS-bounding hypercube ansatz can be efficiently obtained. Our POPS-constrained ansatz are shown to produce excellent parameter distributions which give accurate prediction of test error distributions and highly informative worst-case bounds on model predictions, in the important misspecified, near-deterministic regime where existing methods fail.
Synthetic test data
All synthetic data was generated with Python/JAX code, to be made available following peer review.
Atomic simulation data
Atomic training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022). 10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components). Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), which can be expressed as a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
@@ Acknowledgements
This work was initiated during a visit at the Institute for Pure and Applied Mathematics at the University of California, Los Angeles (supported by National Science Foundation (NSF) grant DMS-1925919) and continued at the Institute for Mathematical and Statistical Innovation, University of Chicago (supported by the NSF grant DMS-1929348). Their hospitality is graciously acknowledged. TDS gratefully acknowledges support from ANR grants ANR-19-CE46-0006-1 and ANR-23-CE46-0006-1, IDRIS allocation A0120913455 and an Emergence@INP grant from the CNRS. D.P. was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20220063DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).
@@ Appendix A PAC-Bayes concentration inequalities for regression
Determining the value of the upper bound for $C_{0}[\pi_{0}]$ in the PAC-Bayes concentration inequality (21) requires certain assumptions on the distribution of the error between the empirical and true expected loss (Germain et al., 2016). Following (Germain et al., 2016), we assume that this error distribution is unbounded but sub-normal, i.e. all error fluctuations are bounded by a normal distribution of variance $s^{2}$, giving $C_{0}[\pi_{0}]<s^{2}/2$. We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion. For completeness, we recall here the bound used in the main text:
${\mathcal{L}}[\pi]\leq\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi}{\pi_{0}}\right|{\rm d}\pi({\bm{\Theta}})-\frac{\ln|\xi|}{N}+C_{0},$ (21)
@@ Appendix B Model selection criteria for deterministic models
We note that the value of the PAC-Bayes loss bound (21) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in (Germain et al., 2016). In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected. However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed. Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{E}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (23)
@@ Appendix C Summary of Laplace’s method
Appendix D POPS-constrained ensemble ansatz
Selecting some set of pointwise fits and $\mathcal{T}_{E}=\{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\}_{\mathcal{X}}$ and weights $\mathcal{W}_{E}=\{{\rm w}_{E}({\mathbf{X}})>0\}_{\mathcal{X}}$ such that $\langle{\rm w}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$, the covering constraint is satisfied by
[…]
For any choice of weights $\mathcal{W}_{E}$, an approximate optimal set ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$ can be determined in a variational setting at finite $N$, using (25) in the PAC-Bayes loss bound (21) for the loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$. As $P/N\to 0$ we show in the sub-section below this reduces to the set of POPS-constrained loss minimizers
${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E},\quad{\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}})$ (26)
The POPS-constrained loss-minimization ensures the fits will be as close as possible under the Fisher metric ${\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}$, such that ${\bm{\Theta}}^{*}_{\mathbf{X}}\to{\bm{\Theta}}^{*}_{\mathcal{L}}$ for specified models. Under misspecification, the generalization error of (25) using $\mathcal{T}^{*}_{E}$ has the $P/N\to 0$ limit
${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}}.$ (27)
In general we expect a uniform weighting ${\rm w}_{E}({\mathbf{X}})=1$ to be sub-optimal for generalization as the loss-minimizing fits ${\bm{\Theta}}^{*}_{\mathbf{X}}$ will be too close to ${\bm{\Theta}}^{*}_{\mathcal{L}}$.
To optimize ${\rm w}_{E}({\mathbf{X}})$ we define a density $\rho_{{\mathbf{X}}}^{*}({\mathbf{X}}^{\prime})={\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})/\langle{\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}$ of pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}}$ satisfying $|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}})-{\mathcal{E}}({\mathbf{X}})|=\mathcal{O}(\epsilon)$, then use Jensen’s inequality to write the ”loss-like” bound
${\mathcal{G}}_{0}[\pi_{E}]$ $={\mathcal{G}}_{U}-\left\langle\ln\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime})\rho_{{\mathbf{X}}}^{*}({\mathbf{X}}^{\prime})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}}$
$\leq{\mathcal{G}}_{U}-\left\langle\ln|{\rm w}_{E}({\mathbf{X}}^{\prime})|\left\langle\rho_{{\mathbf{X}}}^{*}({\mathbf{X}}^{\prime})\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}},$ (28)
where ${\mathcal{G}}_{U}$ is the generalization error (27) with uniform weighting ${\rm w}_{E}({\mathbf{X}})=1$. Under the constraint $\langle{\rm w}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$ it is simple to show this upper bound is minimized by
${\rm w}^{*}_{E}({\mathbf{X}})\in\mathcal{W}_{E}^{*},\quad{\rm w}^{*}_{E}({\mathbf{X}})=\lambda\left\langle\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}.$ (29)
where $\lambda$ ensures $\langle{\rm w}^{*}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$. Evaluation requires at most $\mathcal{O}(N^{2})$ effort but only $\mathcal{O}(N)$ for storage and resampling. The final ensemble ansatz parameter posterior then reads
$\pi^{*}_{E}({\bm{\Theta}})\equiv\langle{\rm w}^{*}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}..$ (30)
Equation (30) is a variationally optimized ensemble ansatz for misspecified regression problems. The worst case of extreme misspecification, corresponding to non-intersecting POPS ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}^{\prime})\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$, gives $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})=\delta_{{\mathbf{X}}{\mathbf{X}}^{\prime}}$ and thus uniform weights ${\rm w}^{*}_{E}({\mathbf{X}})=1$. Interestingly, it is simple to see we have uniform weighting at large $\epsilon$ as $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})$ is uniform. For real systems we find ${\rm w}^{*}_{E}({\mathbf{X}})$ is highly non-uniform and smallest for points near to ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as expected. However, numerical test reveal moments of the predicted ensemble errors are still underestimated, as the ”loss like” bounds still bring all points too close to the minimum loss value, which motivates the hypercube approach described in the text. However, the max/min bounds of the ensemble, which are insensitive to our choice of ${\rm w}({\bf X})$, remain robust, as shown in figure 1.
D.1 Variational optimization